OpenStudy (anonymous):

Solve S = 4v2 for v.

OpenStudy (anonymous):

is that S=4v^2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok well lets start by geting v^2 by itself

OpenStudy (anonymous):

ok S/4 = v^2

OpenStudy (anonymous):

yup now sqrt the entire problem

OpenStudy (anonymous):

\[\sqrt{s/4} = v\]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

hope that helped